Hydrogen peroxide solution (20 mL) reacts quantitatively with a solution of KMnO4 (20 mL) acidified with dilute H2SO4. The same volume of the KMnO4 solution is just decolourised by 10 mL of MnSO4 in neutral medium simultaneously forming a dark brown precipitate of hydrated MnO2. The brown precipitate is dissolved in 10 mL of 0.2 M sodium oxalate under boiling condition in the presence of dilute H2SO4. Write the balanced equations involved in the reactions and calculate the molarity of H2O2.

Grade:11

1 Answers

Navjyot Kalra
9 years ago
Sol. Meq of oxalate = 10 × 0.2 × 2 = 4 Meq of MnO2 formed = Meq of oxalate = 4 Meq of KMnO4 in 20 mL = 4 ⇒ Normality of H2O2 × 20 = 4 ⇒ Normality of H2O2 = 0.20 N ⇒ Molarity of H2O2 = 0.20/2 = 0.10 M The balanced reactions are : 2KMnO4 + 5H2O2 + 3H2SO4 → 2MnSO4 + 502 + K2SO4 + 8H2O MnO2 + Na2C2O4 + 2H2SO4 → MnSO4 + Na2SO4 + 2CO2 + 2H2O